Two Masses, m and 2m, approach each other along a path at right angles to each other . After collision, they stick together and move off at 2m/s at angle 37⁰ to the original direction of the mass m. What were the initial speeds of the two particles ??
1
Expert's answer
2021-08-24T06:53:04-0400

Let us denote initial velocities of the masses as "v_1" and "v_2". The law of conservation of a momentum gives:

"m{\bf v_1}+2m{\bf v_2}=(m+2m)\bf u"

In projection:


"mv_1=3mu\cos 37^{\circ}\\
2mv_2=3mu\sin 37^{\circ}"

Hence,

"v_1=3*2\:{\rm m /s}\cos37^{\circ}=4.8\:{\rm m/s}\\
v_2=3/2*2\:{\rm m /s}\sin37^{\circ}=1.8\:{\rm m/s}"